OpenStudy (anonymous):

If f(x)= 2x 2 + 3, then f(x - 1)= 2x 2 + 4x + 5. is this correct?

OpenStudy (ivettef365):

the problem is not well defined, please clarify the equation of f(x)= ?? what is 2x2

OpenStudy (anonymous):

\[f \left( x-1 \right)=2\left( x-1 \right)^{2}+3\] solve and get the solution.